OpenStudy (anonymous):

Differential Equation: y' = y-y^3. I've got it down to: y^2 - 2y = e^(2t) e^c But the problem is: List all constant solutions for the DE I don't quite get how to get any further. x_x

OpenStudy (jamesj):

If y' = y - y^3 then \[ \frac{dy}{y(1-y^2)} = dx \] Now integrate and you'll find the expression you've written down isn't quite right.

OpenStudy (anonymous):

oh :c Lemme redo it. @_@ What does a 'constant solution' mean, though? Is that the same thing as equilibrium solution?

OpenStudy (anonymous):

I get a really weird solution full of radicals...

OpenStudy (anonymous):

Constant solutions are the same as the equilibrium solutions

OpenStudy (anonymous):

*facepalm* Constant solution is equilibrium solution. I'm stupid and didn't read the problem right. didn't even need to do the actual integration Q_Q Sorries. I have the answer now :3

OpenStudy (jamesj):

Ah, well, a constant is a solution if y' = 0. I.e., y - y^3 = 0 => y = 0, 1 or -1.

OpenStudy (anonymous):

:) be glad you don't have to, that dq is real dirty

OpenStudy (anonymous):

Thank you~ That's exactly what I got O:

OpenStudy (anonymous):

@Fronklin Yea ;_; I forgot how to do integration by parts aha...